# MLPs Learn In-Context

### v1

@@ Abstract
In-context learning (ICL), the remarkable ability to solve a task from only input exemplars, has commonly been assumed to be a unique hallmark of Transformer models. In this study, we demonstrate that multi-layer perceptrons (MLPs) can also learn in-context. Moreover, we find that MLPs, and the closely related MLP-Mixer models, learn in-context competitively with Transformers given the same compute budget. We further show that MLPs outperform Transformers on a subset of ICL tasks designed to test relational reasoning. These results suggest that in-context learning is not exclusive to Transformers and highlight the potential of exploring this phenomenon beyond attention-based architectures. In addition, MLPs’ surprising success on relational tasks challenges prior assumptions about simple connectionist models. Altogether, our results endorse the broad trend that “less inductive bias is better" and contribute to the growing interest in all-MLP alternatives to task-specific architectures.
School of Engineering and Applied Sciences,
Center for Brain Science,
Kempner Institute for the Study of Natural and Artificial Intelligence
Harvard University, Cambridge, MA 02138
@@ 1 Introduction
The last few years have witnessed meteoric progress in neural language models. Catalyzed by the Transformer architecture and driven by a steady increase in scale, these aptly-named Large Language Models (LLMs) demonstrate unprecedented competence in drafting grammatical text, answering questions, summarizing content, generating creative output, and even reasoning through non-trivial puzzles [1, 2, 3].
Crucial to an LLM’s proficiency is its ability to learn in-context [4, 5, 2]. In-context learning (ICL) refers to a task paradigm where exemplars from a novel task are presented during inference time rather than during training (Figure 1a). The model must then respond correctly to a query based only on these novel exemplars. No weight updates occur throughout this process; rather, the model infers the task from the input exemplars and, despite having fixed weights, produces the correct output.
ICL is commonly assumed to be a unique ability of Transformers, and explanations of the phenomenon often ground their constructions in attention-based architectures [6, 7, 8, 9, 10]. In contrast, we demonstrate that the simple multi-layer perceptron (MLP) can also learn in-context — and moreover, learn in-context competitively with the full Transformer given the same compute budget. [^1] Universal approximation by MLPs suggests that they may be able to learn in-context, though at uncertain cost. We demonstrate that MLPs learn in-context without significantly larger compute (and occasionally quite a bit smaller) than Transformers. These results suggest that ICL is not exclusive to Transformers, and highlights the potential for studying the phenomenon beyond attention-based architectures. Additionally, the emerging narrative that “less inductive bias is better" [11, 12, 13] suggests that as data and compute resources grow, models with weaker inductive bias like MLPs will outperform models with stronger inductive bias like Transformers. Equipped with architectural constraints like self attention and positional encodings, Transformers may be more strongly biased towards solutions that are less aligned with certain task structures than those discovered by the more general MLP [12]. We support this heuristic by observing how MLPs outperform Transformers on a subset of ICL tasks designed to probe relational reasoning, as well as on simple (non-ICL) regression and classification (in Appendix B).
@@
In comparing architectures, we emphasize controlled, representative tasks commonly studied in the literature as testbeds for ICL, where the specific capacity for in-context learning can be precisely characterized. In Section 2, we examine ICL versions of regression [14, 6, 15, 8] and classification [9, 16]. As the two primary task paradigms of machine learning, regression and classification form a representative basis in comparing ICL capacity. In Section 3, we consider a subset of ICL classification tasks used in the cognitive science literature to probe relational reasoning [17, 18, 19]. These tasks are nontrivial ICL examples where MLPs outperform Transformers, and also challenge common beliefs about MLPs’ proficiency at relational reasoning (additional details in Appendix A). Overall, by focusing on controlled tasks, we avoid possible confounds irrelevant to ICL introduced by complex settings like vision and natural language, though our overall conclusions are consistent with existing results in these domains [12, 20, 21, 11].
@@
To ensure different architectures are comparable across tasks, we observe the following ground rules. First, we compare models based on the total compute required for training (measured in peta-floating point operations, PFLOPs), which summarizes influences like parameter count, training iterations, and architectural efficiency. Details on how we compute this quantity are provided in Appendix C.13. Measuring by compute reflects the practical use of these models, fairly compares architectures by performance per floating-point cost, and is an established scale for describing neural scaling laws [22]. Second, where a single model is required, we use the largest tested parameter setting, keeping compute cost equal across architectures by varying training iterations. Doing so ensures that the results are not bottlenecked by model size. As shown in Appendix D, selecting smaller models may occasionally result in suboptimal performance plateaus, even if training iterations are increased to match the compute of larger models. Specific model and task configurations are enumerated in Appendix C.
@@ 1.1 Related work
In-context learning has been studied in a number of controlled settings. In particular, ICL has been reproduced for linear regression, where a Transformer trained to perform the task can extrapolate to novel input/label pairs provided in-context [14, 6, 15, 23, 24, 25, 10]. Proposed mechanisms whereby a Transformer accomplishes the feat include that the Transformer implements some form of gradient descent [7, 6] or recapitulates least-squares or Ridge regression [8, 6, 10]. It has also been observed that a Transformer interpolates between in-weight learning (IWL), the traditional paradigm where the model learns specific examples through training, to in-context learning, where the model uses only exemplars provided in the input context at inference time [15, 23]. Such a transition occurs as a function of data diversity, where datasets with richer variety encourage the development of ICL competency. Analogous phenomena have been observed in in-context classification tasks [16, 9]. Impressively, the ICL performance attained in these tasks by Transformers approaches Bayes optimality [26, 24, 25, 27, 10].
@@
These studies nearly all ground their investigations in Transformer models, and explicitly assume that the model uses an attention mechanism to implement ICL. The exceptions include Chan et al. [16], who discover that recurrent neural networks (both vanilla RNNs and LSTMs) are unable to learn an in-context classification task under the same conditions where a Transformer can, and Xie et al. [26], who discover that LSTMs can in fact learn in-context on a synthetic language modeling task. Recently, Lee et al. [28] found that a wide variety of causal sequence models can learn in-context on a broad array of toy tasks, with varying degrees of success. Park et al. [29] support this finding by showing how state space models and their hybrids with Transformers can learn in-context competitively. To the best of our knowledge, no prior work has examined in-context learning in vanilla MLPs.
@@
The current resurgence of interest in applying MLPs to modern, complex tasks originates with Tolstikhin et al. [12], which introduced the MLP-Mixer model. Mixers operate by alternating MLPs across the dimensions of the input, treating the remaining dimensions as batch dimensions. Despite their simplicity, Mixers attain state-of-the-art performance on image classification, recalling the folk heuristic that “less inductive bias is better" [13, 11]. In the ensuing years, “all-MLP" models based primarily on MLP components have exploded across many areas including vision [11] and natural language [20, 21]. While strong performance has been documented on natural language, less is known about MLPs’ specific proficiency for ICL, and how it compares with Transformer models. In this study, we select a series of controlled, representative tasks that clarify an MLP’s surprising competence for ICL. Our findings underscore the ultimate utility of MLPs, uncovering avenues of both theoretic and practical interest.
@@ 2.1 ICL regression
We present in-context regression as follows [14, 8]. The input consists of a sequence of values $(\boldsymbol{x}_{1},y_{1}),(\boldsymbol{x}_{2},y_{2}),\ldots,(\boldsymbol{x}_{L},y_{L})$, where $\boldsymbol{x}_{i}\in\mathbb{R}^{n}$ and $y_{i}\in\mathbb{R}$. The $\boldsymbol{x}_{i},y_{i}$ pairs are linearly related through a set of weights $\boldsymbol{\beta}\in\mathbb{R}^{n}$ such that $y_{i}=\boldsymbol{x}_{i}\cdot\boldsymbol{\beta}+\varepsilon$, with noise $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$. Finally, the input includes a query $\boldsymbol{x}_{q}$. The model output is a single scalar regressed against the corresponding $y_{q}$. Crucially, the weights $\boldsymbol{\beta}$ vary between input sequences. The model cannot rely on learning any one $\boldsymbol{\beta}$. Rather, it must infer from context exemplars $(\boldsymbol{x}_{i},y_{i})$ what the corresponding $\boldsymbol{\beta}$ must be, and use this to predict the correct output $y_{q}$. Figure 1b illustrates the task, with additional details in Appendix C.
Following Raventós et al. [15], we consider two different task distributions: finite and unrestricted. For the finite distribution, we fix a finite pool of weights before training $\boldsymbol{\beta}_{1},\boldsymbol{\beta}_{2},\ldots,\boldsymbol{\beta}_{k}$, where $\boldsymbol{\beta}_{i}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}/n)$. For each input, we sample a new $\boldsymbol{\beta}$ by selecting uniformly at random one weight from the pool $\left\{\boldsymbol{\beta_{i}}\right\}_{i=1}^{k}$. Larger $k$ corresponds to higher data diversity. For the unrestricted distribution, a new set of weights is sampled for each input $\boldsymbol{\beta}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}/n)$. The unrestricted distribution can be thought of as the $k\rightarrow\infty$ case, and requires full ICL competency in order to infer the correct weights relating the context exemplars. Unless otherwise stated, we use $n=8$ dimensional inputs.
@@ Results.
One final observation: as data diversity increases, Transformers exhibit a transition from in-weight learning (IWL), the traditional paradigm where the model learns specific examples through training, to in-context learning, where the model uses only context exemplars presented at inference time [15]. We next show that MLPs exhibit a similar transition. Following Raventós et al. [15], we train each model on a finite distribution with $k$ fixed regression weights. As we increase $k$, we record the MSE obtained by each model on both the finite distribution $\boldsymbol{\beta}\sim\mathcal{U}\left(\left\{\boldsymbol{\beta_{i}}\right\}_{i=1}^{k}\right)$ using the same $\boldsymbol{\beta}$’s from training (Figure 1e) and the unrestricted distribution $\boldsymbol{\beta}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}/n)$ where $\boldsymbol{\beta}$’s are sampled anew (Figure 1f). We determine whether a model has learned the in-weight solution by comparing its MSE to that of the discrete minimum mean squared error (dMMSE) estimator, which is a Bayesian estimator derived from a prior matched to the finite training distribution (see Appendix C.6 for details).  [^2] The dMMSE estimator averages across the $k$ weights in the finite training distribution based on their fit to the current context exemplars. We characterize the in-context solution by a Ridge estimator with the Bayes optimal choice of regularization. For small $k$, all models demonstrate in-weight learning by tracing the dMMSE curve. As $k$ increases, we observe a swift transition to the Ridge curve, indicating a transition to in-context learning. The Transformer makes this transition at a somewhat smaller $k$ than the MLP models. We consider additional plots and parameterizations in Appendix D.
Figure 1: ICL regression and classification results. (a) ICL presents context exemplars from a novel task (red), followed by a query input (blue). The model must infer the solution (green) based on the context. (b) ICL regression. (c) Compute vs. MSE on the unrestricted task distribution. Each point represents a single model, with particular parameters and training iterations. At large compute, MSE is approximately equal across all architectures. The red line corresponds to the Bayes optimal Ridge MSE. (d) Excess MSE (MSE above Bayes optimal) for varying context length $L$ on the unrestricted task distribution. Excess MSE remains flat for Mixers, but rises somewhat for Transformers. MLPs fail to learn in-context at all beyond $2^{6}$ context exemplars. The grey line corresponds to the excess MSE incurred by always guessing zero. (e, f) IWL to ICL transition with increasing data diversity. We train on a finite distribution with $k$ weights, then test on both the finite training distribution and the unrestricted distribution. All models exhibit a transition from IWL (represented by dMMSE) to ICL (represented by Ridge) as $k$ increases. Note: it is possible to “outperform" Bayes optimal Ridge on the finite training distribution by learning in-weight the underlying $\boldsymbol{\beta}$’s. (g) ICL classification, with burstiness $B=3$. Multiple clusters may share the same label. (h) Compute vs. cross entropy loss on ICL classification, with $k=2048$ clusters, $B=4$, and $L=8$, which pushes all models to learn in-context. At large compute, all architectures attain near-zero cross entropy loss. The gray line corresponds to loss obtained from placing equal probability on the 2 (of $C=32$) labels present in context. (i) Cross entropy loss for varying context length $L$ on the task configuration in (h). Loss is relatively flat for all architectures, though it increases a little for Mixers. (j) IWL to ICL transition with increasing data diversity, where $L=8$ and $B=4$. All models exhibit a transition from IWL to ICL as the number of clusters $k$ increases. (all) We use $n=8$ dimension inputs. All line plots feature 95 percent confidence intervals about the mean, estimated from 5 replications.

### v3

@@ Abstract
In-context learning (ICL), the remarkable ability to solve a task from only input exemplars, is often assumed to be a unique hallmark of Transformer models. By examining commonly employed synthetic ICL tasks, we demonstrate that multi-layer perceptrons (MLPs) can also learn in-context. Moreover, MLPs, and the closely related MLP-Mixer models, learn in-context comparably with Transformers under the same compute budget in this setting. We further show that MLPs outperform Transformers on a series of classical tasks from psychology designed to test relational reasoning, which are closely related to in-context classification. These results underscore a need for studying in-context learning beyond attention-based architectures, while also challenging prior arguments against MLPs’ ability to solve relational tasks. Altogether, our results highlight the unexpected competence of MLPs in a synthetic setting, and support the growing interest in all-MLP alternatives to Transformer architectures. It remains unclear how MLPs perform against Transformers at scale on real-world tasks, and where a performance gap may originate. We encourage further exploration of these architectures in more complex settings to better understand the potential comparative advantage of attention-based schemes.
@@ 1 Introduction
The last few years have witnessed meteoric progress in neural language models. Catalyzed by the Transformer architecture and driven by a steady increase in scale, these aptly-named Large Language Models (LLMs) demonstrate unprecedented competence in drafting grammatical text, answering questions, summarizing content, generating creative output, and even reasoning through non-trivial puzzles (Bubeck et al., 2023; Brown et al., 2020; Achiam et al., 2023).
Crucial to an LLM’s proficiency is its ability to learn in-context (Lu et al., 2023; Dong et al., 2022; Brown et al., 2020). In-context learning (ICL) refers to a task paradigm where exemplars from a novel task are presented during inference time rather than during training (Figure 1a). The model must then respond correctly to a query based only on these novel exemplars. No weight updates occur throughout this process; rather, the model infers the task from the input exemplars and, despite having fixed weights, produces the correct output.
ICL is commonly assumed to be a unique ability of Transformers, and explanations of the phenomenon often ground their constructions in attention-based architectures (Akyürek et al., 2024; Von Oswald et al., 2023; Zhang et al., 2023; Reddy, 2024; Lu et al., 2024). On controlled regression and classification tasks targeted specifically for evaluating ICL, we demonstrate that the simple multi-layer perceptron (MLP) can also learn in-context — and moreover, learn in-context competitively with the full Transformer given the same compute budget. [^1] Universal approximation by MLPs suggests that they may be able to learn in-context, though at uncertain cost. We demonstrate that MLPs learn in-context without significantly larger compute (and occasionally quite a bit smaller) than Transformers. These results suggest that ICL is not an exclusive feature of attention-based architectures, and highlights the need for studying the phenomenon in a broader setting.
@@
We focus on controlled tasks commonly studied in the ICL literature, where the specific capacity for in-context learning can be precisely characterized. These tasks are necessarily synthetic approximations of natural language ICL prompting, but allow us to disambiguate a model’s capacity for in-context learning from its ability to attain natural language fluency. In Section 2, we examine ICL versions of regression (Garg et al., 2022; Akyürek et al., 2024; Raventós et al., 2024; Zhang et al., 2023) and classification (Reddy, 2024; Chan et al., 2022). As the two primary task paradigms of machine learning, regression and classification form a representative basis for measuring ICL competency. In Section 3, we consider a series of classical tasks used in the psychology literature to probe relational reasoning (Campbell et al., 2023; Skinner, 1950; Sablé-Meyer et al., 2021), which are functionally in-context classification. On these tasks, we find that MLPs outperform Transformers, challenging common beliefs about MLPs’ proficiency at relational reasoning (see Appendix A for an extended discussion). In focusing on controlled tasks, we avoid confounds irrelevant to ICL introduced by naturalistic settings like language and vision. Nonetheless, our findings remain consistent with existing results that do test MLPs in these more complex domains (Tolstikhin et al., 2021; Liu et al., 2021; Fusco et al., 2022; Bachmann et al., 2024).
@@
To ensure different architectures are comparable across tasks, we observe the following ground rules. First, we compare models based on the total compute required for training (measured in peta-floating point operations, PFLOPs), which summarizes influences like parameter count, training iterations, and architectural efficiency. Details on how we compute this quantity are provided in Appendix C.13. Measuring by compute reflects the practical use of these models, fairly compares architectures by performance per floating-point cost, and is an established scale for defining neural scaling laws (Kaplan et al., 2020). Second, where a single model is required, we select the best model configuration as measured by loss, keeping compute cost equal across architectures. Data are presented online, reflecting the “one-pass" setting common in training large language models (Brown et al., 2020). Specific model and task configurations are enumerated in Appendix C.
@@ 1.1 Related work
In-context learning has been widely studied in a number of controlled settings. In particular, ICL has been reproduced for linear regression, where a Transformer trained to perform the task can extrapolate to novel input/label pairs provided in-context (Garg et al., 2022; Akyürek et al., 2024; Raventós et al., 2024; Wu et al., 2024; Bai et al., 2024; Li et al., 2023; Lu et al., 2024). Proposed mechanisms whereby a Transformer accomplishes the feat include that the Transformer implements some form of gradient descent (Von Oswald et al., 2023; Akyürek et al., 2024) or recapitulates least-squares or Ridge regression (Zhang et al., 2023; Akyürek et al., 2024; Lu et al., 2024). It has also been observed that a Transformer interpolates between in-weight learning (IWL), the traditional paradigm where the model learns specific examples through training, to in-context learning, where the model uses only exemplars provided in the input context at inference time (Raventós et al., 2024; Wu et al., 2024). Such a transition occurs as a function of data diversity, where datasets with more distinct examples encourage the development of ICL competency. Analogous phenomena have been observed in in-context classification tasks (Chan et al., 2022; Reddy, 2024). Impressively, the ICL performance attained in these tasks by Transformers approaches Bayes optimality (Xie et al., 2021; Bai et al., 2024; Li et al., 2023; Ahuja et al., 2024; Lu et al., 2024).
@@
These studies nearly all ground their investigations in Transformer models, and explicitly assume that the model uses an attention mechanism to implement ICL. The exceptions include Chan et al. (2022), who discover that recurrent neural networks (both vanilla RNNs and LSTMs) are unable to learn an in-context classification task under the same conditions where a Transformer can, and Xie et al. (2021), who discover that LSTMs can in fact learn in-context on a synthetic language modeling task. Recently, Lee et al. (2024) found that a wide variety of causal sequence models can learn in-context on a broad array of toy tasks, with varying degrees of success. Park et al. (2024) support this finding by showing how state space models and their hybrids with Transformers can learn in-context competitively. To the best of our knowledge, no prior work has examined in-context learning in vanilla MLPs.
@@
The current resurgence of interest in applying MLPs to modern, complex tasks originates with Tolstikhin et al. (2021), which introduced the MLP-Mixer model. Mixers operate by alternating MLPs across the dimensions of the input, treating the remaining dimensions as batch dimensions. Despite their simplicity, Mixers attain state-of-the-art performance on image classification, recalling the broad observation that “less inductive bias is better" (Sutton, 2019; Bachmann et al., 2024). In the ensuing years, “all-MLP" models based primarily on MLP components have spread across many areas including vision (Bachmann et al., 2024) and natural language (Liu et al., 2021; Fusco et al., 2022). While strong performance has been documented on natural language, less is known about MLPs’ specific proficiency for ICL, and how it compares with Transformer models. In this study, we select a series of controlled, representative tasks that clarify an MLP’s surprising competence for ICL. Our findings underscore the ultimate utility of MLPs, uncovering avenues of both theoretic and practical interest.
Figure 1: ICL regression and classification results. (a) ICL presents context exemplars from a novel task (red), followed by a query input (blue). The model must infer the solution (green) based on the context. (b) ICL regression example. The model receives linearly-related input points, and must regress the query point. (c) Compute vs. MSE on the unrestricted task distribution. Each point represents a single model, with particular parameters and training iterations. At large compute, MSE is approximately equal across all architectures. The red line corresponds to the Bayes optimal Ridge MSE. (d) Excess MSE (MSE above Bayes optimal) for varying context length $L$ on the unrestricted task distribution. Excess MSE remains flat for Mixers, but rises somewhat for Transformers. MLPs fail to learn in-context at all beyond $2^{6}$ context exemplars. The grey line corresponds to the excess MSE incurred by always guessing zero. (e, f) IWL to ICL transition with increasing data diversity. We train on a finite distribution with $k$ weights, then test on both the finite training distribution and the unrestricted distribution. All models exhibit a transition from IWL (represented by dMMSE) to ICL (represented by Ridge) as $k$ increases. Note: it is possible to “outperform" Bayes optimal Ridge on the finite training distribution by learning in-weight the underlying $\boldsymbol{\beta}$’s. (g) ICL classification example, with burstiness $B=3$. Multiple clusters may share the same label. (h) Compute vs. cross entropy loss on ICL classification, with $k=2048$ clusters, $B=4$, and $L=8$, which pushes all models to learn in-context. At large compute, all architectures attain near-zero cross entropy loss. The gray line corresponds to loss obtained from placing equal probability on the 2 (of $C=32$) labels present in context. (i) Cross entropy loss for varying context length $L$ on the task configuration in (h). Loss is relatively flat for all architectures, though it increases a little for Mixers. (j) IWL to ICL transition with increasing data diversity, where $L=8$ and $B=4$. All models exhibit a transition from IWL to ICL as the number of clusters $k$ increases. (all) We use $n=8$ dimension inputs. All line plots feature 95 percent confidence intervals about the mean, estimated from 5 replications.
@@ 2.1 ICL regression
We present in-context regression following its common formulation (Garg et al., 2022; Zhang et al., 2023). The input consists of a sequence of values $(\boldsymbol{x}_{1},y_{1}),(\boldsymbol{x}_{2},y_{2}),\ldots,(\boldsymbol{x}_{L},y_{L})$, where $\boldsymbol{x}_{i}\in\mathbb{R}^{n}$ and $y_{i}\in\mathbb{R}$. The $\boldsymbol{x}_{i},y_{i}$ pairs are linearly related through a set of weights $\boldsymbol{\beta}\in\mathbb{R}^{n}$ such that $y_{i}=\boldsymbol{x}_{i}\cdot\boldsymbol{\beta}+\varepsilon$, with noise $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$. Finally, the input includes a query $\boldsymbol{x}_{q}$. The model output is a single scalar regressed against the corresponding $y_{q}$. Crucially, the weights $\boldsymbol{\beta}$ vary between input sequences. The model cannot rely on learning any one $\boldsymbol{\beta}$. Rather, it must infer from context exemplars $(\boldsymbol{x}_{i},y_{i})$ what the corresponding $\boldsymbol{\beta}$ must be, and use this to predict the correct output $y_{q}$. Figure 1b illustrates the task, with additional details in Appendix C.
In the main text, our task fixes the number of context points at $L$. A common variation on this tasks allows the number of context points to vary, and trains the model autoregressively. Results on this autoregressive variation are presented in Figure 5, and are unchanged from the fixed-length case.
Following Raventós et al. (2024), we consider two different task distributions: finite and unrestricted. For the finite distribution, we fix a finite pool of weights before training $\boldsymbol{\beta}_{1},\boldsymbol{\beta}_{2},\ldots,\boldsymbol{\beta}_{k}$, where $\boldsymbol{\beta}_{i}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}/n)$. For each input, we sample a new $\boldsymbol{\beta}$ by selecting uniformly at random one weight from the pool $\left\{\boldsymbol{\beta_{i}}\right\}_{i=1}^{k}$. Larger $k$ corresponds to higher data diversity. For the unrestricted distribution, a new set of weights is sampled for each input $\boldsymbol{\beta}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}/n)$. The unrestricted distribution can be thought of as the $k\rightarrow\infty$ case, and requires full ICL competency in order to infer the correct weights relating the context exemplars. Unless otherwise stated, we use $n=8$ dimensional inputs.
@@ Results.
One final observation: as data diversity increases, Transformers exhibit a transition from in-weight learning (IWL), the traditional paradigm where the model learns specific examples through training, to in-context learning, where the model uses only context exemplars presented at inference time (Raventós et al., 2024). We next show that MLPs exhibit a similar transition. Following Raventós et al. (2024), we train each model on a finite distribution with $k$ fixed regression weights. As we increase $k$, we record the MSE obtained by each model on both the finite distribution $\boldsymbol{\beta}\sim\mathcal{U}\left(\left\{\boldsymbol{\beta_{i}}\right\}_{i=1}^{k}\right)$ using the same $\boldsymbol{\beta}$’s from training (Figure 1e) and the unrestricted distribution $\boldsymbol{\beta}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}/n)$ where $\boldsymbol{\beta}$’s are sampled anew (Figure 1f). We determine whether a model has learned the in-weight solution by comparing its MSE to that of the discrete minimum mean squared error (dMMSE) estimator, which is a Bayesian estimator derived from a prior matched to the finite training distribution (see Appendix C.6 for details). [^2] The dMMSE estimator averages across the $k$ weights in the finite training distribution based on their fit to the current context exemplars. We characterize the in-context solution by a Ridge estimator with the Bayes optimal choice of regularization. For small $k$, all models demonstrate in-weight learning by tracing the dMMSE curve. As $k$ increases, we observe a swift transition to the Ridge curve, indicating a transition to in-context learning. The Transformer makes this transition at a somewhat smaller $k$ than the MLP models. We consider additional plots and parameterizations in Appendix D.